Suppose that you own 2,500 shares of Nocash Corp. and the company is about to pay a 25% stock dividend. The stock currently sells at $100 per share.

a. What will be the number of shares that you hold after the stock dividend is paid? (Do not round intermediate calculations.)

b. What will be the total value of your equity position after the stock dividend is paid? (Do not round intermediate calculations.)

c. What will be the number of shares that you hold if the firm splits five-for-four instead of paying the stock dividend?
Answer #1

a.

Number of Shares before Stock Dividend = 2,500

Number of shares after Stock Dividend = 1.25(2,500) = 3,125 shares

b.

Total Value of Equity after Stock Dividend = 3,125(100) = $312,500

c.

Number of Shares after 5 for 4 split = 2,500(5/4)

Number of Shares after 5 for 4 split = 3125